A Restaurant is open only for 25 days in a month. Expenses for the restaurant include raw material for each sandwich at $5.00 per slice, $1,028.00 as monthly rental and $372.00 monthly as insurance. They consider the cost of lost sales as $4.00 per item. They are able to sell any leftover sandwiches for $3. They prepares 200.00 sandwiches and sells them at a rate of $15.00/sandwich. Today there was a party at nearby office so the demand for sandwiches rose to 228.00. How much profit did the restaurant earn today
Answer #1

Fixed cost per month = rental + insurance = $1028 + $372 = 1400

Fixed cost per working day = 1400/ 25 = $56

Variable cost per sandwich = $5

They prepared 200 sandwiches today. Demand for today = 228. Hence lost sales = 28.

Cost of lost sales = 28*4 = $112

Hence total cost = (200* 5 )+ 56 + 112 = $1168

Total Revenue = Number of sandwiches sold * selling price per item = 200*15 = $ 3000

Hence profit = $3000 - $1168 = $1832